# Federated Instruction Tuning of LLMs with Domain Coverage Augmentation

### v4

@@ A.3.2 Direct Retrieval
A.4 Projector
@@
A.5 Train and Test Dataset Information
Here we provide the train and test dataset details of code (CodeAlpaca [^2] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k), medical (MedAlpaca [^3] https://huggingface.co/datasets/medalpaca), financial (FinGPT (Yang et al., 2023a)) and mathematical (MathInstruct [^4] https://huggingface.co/datasets/TIGER-Lab/MathInstruct) domain, respectively.
Table 8: Train set information of each domain.
Dataset name Domain $N_{sample}$
CodeAlpaca Code 20,022
MedAlpaca Medical 33,955
FinGPT Financial 76,772
MathInstruct Mathematical 224,567
Alpaca General 52,002
As shown in Table 8, the public data consists of four domain-specific instruction datasets and a general instruction dataset, which are CodeAlpaca, MedAlpaca, FinGPT, MathInstruct, and Alpaca, respectively. For each domain’s FedDIT, we randomly select 1000 samples from the in-domain instruction and split them into 10 shards as each client’s local dataset. The rest of the instructions are used as the public dataset $D^{p}$.
Table 9: Test set information of each domain.
Dataset name Domain $N_{sample}$ Metric
HumanEval Code 164 Pass@1
MMLU-Med Medical 1,089 Acc
FPB Financial 152 Acc
FiQA Financial 35 Acc
TFNS Financial 299 Acc
GSM8K Mathematical 1,319 Exact Match
Table 9 shows the detail of the test set used for evaluation: HumanEval is used for the code domain; MMLU-Med is used for the medical domain; FPB, FiQA and TFNS are used for the financial domain, and GSM8K is used for the mathematical domain. Specifically, MMLU-Med uses subjects anatomy, clinical_knowledge, college_biology, college_medicine, medical_genetics and professional_medicine in MMLU.
A.6 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^5] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means (Wu, 2012), we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning (Khosla et al., 2020). We utilize bge-large-en-v1.5666https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2777https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
A.7 Prompts used in the Self-Instruct data generation
@@
A.8 Prompt used for memory extraction attack
@@
A.9 Further Analysis
To further study the effect of different hyperparameters of FedDCA, we undertake a thorough analysis including various retrieval amounts and different k-means cluster numbers $\xi$. In addition, we perform the ablation study on whether using the similarity threshold $\alpha$ in the greedy client center selection.
Effect of Retrieval Number. We report the performance and the corresponding domain coverage of FedDCA with different retrieval amounts on the four domains in Figure 11, respectively. We can see that the domain coverage of FedDCA is increasing along with the retrieval amount in different trends, as well as the performance. Specifically, the domain coverage of each domain increases by 6.36%, 18.69%, 5.04%, and 8.06% in relative, respectively. Along with domain coverage increasing, the performance of FedDCA is increasing by 3.05%, 2.20%, 4.08%, and 6.95% for each domain. Noted that although the code domain is more about following a certain paradigm, which could perform well with a few data and more fine-tuning rounds, it still could benefit from the instruction augmentation.
Figure 11: Effect of different retrieval amounts on the performance of FedDCA and its domain coverage. We show the results on four domains separately. Here, we use the FPB test set to evaluate the performance in the financial domain.
Impact of Different Cluster Number. The hyperparameter $\xi$ is the number of clusters in the k-means algorithm. The experiment is conducted on $\xi=[N,2N,4N,8N]$, where $N$ is the number of clients. Following the setting in Appendix A.6, we set $N=10$. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains in Figure 12, respectively. Results show that there is no best $\xi$ for all domains. Specifically, the best $\xi$ of code, medical, financial, and mathematical domains are 80, 80, 40, and 10, respectively.
Figure 12: Impact of different cluster number $\xi$ on the cross-client domain coverage. We report the domain coverage of the augmented dataset via FedDCA with different cluster numbers $\xi$ on the four domains.
Ablation Study. We conduct the experiment with FedDCA w/o similarity threshold $\alpha$ on the four domains based on the FedAvg FL strategy. The performance and the corresponding domain coverage are shown in Table 10, where FedDCA without the similarity threshold $\alpha$ is marked as FedDCA. The result shows that the performance of FedDCA with similarity threshold $\alpha$ is slightly better than FedDCA without using $\alpha$ in code, financial, and mathematical domains, as the similarity scores in the medical domain are relatively lower. We show the similarity score distribution of the four domains in Figure LABEL:fig:similarity_score. For each domain, we plot each similarity score’s distribution of 10 clients. The similarity score is computed between the selected client center and the public data. Then, we show the similarity score distribution using the histogram plot.
Table 10: Ablation study on the performance and domain coverage of FedDCA w/o similarity threshold $\alpha$.
Metric Performance(%) Domain Coverage
FedDCA FedDCA FedDCA FedDCA
H-Eval 35.97 36.58 0.8972 0.9348
M-Med 73.40 73.40 0.9348 0.9348
FPB 66.25 67.24 0.9353 0.9815
FiQA 23.27 35.27
TFNS 69.34 73.32
GSM8K 51.78 52.46 0.9128 0.9320
A.10 Augmentation strategy visualization
@@
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE (van der Maaten & Hinton, 2008), as shown in Figure LABEL:fig:augmentation_strategy_visualization. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 3. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
A.11 Frequently Used Notation
Table 11: Frequently used notation.
NOTATIONS REMARK
$N,C$ Clients number, client set $C=\{c_{1},c_{2},\ldots,c_{N}\}$.

### v5

@@ A.2.2 Direct Retrieval
Appendix B Related Works
Federated Instruction Tuning.
Instruction tuning has been widely applied across various application areas of large language models (LLM), serving as a key technique to enhance the capabilities and controllability of LLM Zhang et al. [2023c]; Wei et al. [2022]. Recently, federated instruction tuning (FedIT) has emerged as an effective strategy for the distributed optimization of LLMs, leveraging federated learning (FL) protocols to improve the handling of privacy-sensitive tasks in real-world scenarios. So far, several FedIT frameworks Ye et al. [2024b, a]; Zhang et al. [2023b] have been established to evaluate the effectiveness of FedIT across multiple datasets, tasks, and FL methods. While these platforms provide a foundation for research, they have not yet introduced more complex federated algorithms and deeply investigate the challenging problems and factors affecting FedIT, which are crucial for advancing this field.
PrivateLoRA Wang et al. [2023b] addresses privacy and efficiency issues by exploiting the low-rank properties of residual activations to reduce communication costs, significantly lowering the communication overhead through collaborative computation between the server and clients while effectively maintaining the privacy of local data. While FewFedPIT Zhang et al. [2024b] focuses on the few-shot learning setting in FedIT, using self-generated data by pre-trained LLMs locally to mitigate the paucity of data and first discussing memory extraction attacks within FedIT.
Non-IID data distribution is a common challenge in FL Karimireddy et al. [2020]; Li et al. [2020]; Sattler et al. [2019]; Li et al. [2019, 2021]. However, the impact of data heterogeneity on FedIT has not been fully explored. Shepherd Zhang et al. [2023b] constructing heterogeneity based on the topic of general instruction datasets. Results demonstrate that, unlike the consensus of traditional FL, for general FedIT, heterogeneity has a positive effect. By aggregating diverse instructions from clients, the diversity increases, thereby enhancing the model’s adaptability to various tasks. However, it just scratches the surface and does not explore issues in FedDIT. To fully understand the impact of data heterogeneity on FedDIT, we conduct a comprehensive analysis in this work and unveil a crucial finding: there is no monotonic correlation between the degree of non-independent and identically distributed (non-iid) and LLM’s performance in the context of FedDIT (Section 3).
Domain Instruction Augmentation.
In the real world, there is an urgent need for training LLMs with specific functionalities (e.g., reasoning capabilities) or domain-specific LLMs (e.g., code Nijkamp et al. [2023]; Luo et al. [2024], medical Zhang et al. [2023d], financial Yang et al. [2023b, a]; Zhang et al. [2023a]; Wu et al. [2023], mathematical Yue et al. [2024]; Luo et al. [2023]). Existing works tend to use open-source domain-specific instruction tuning datasets for training. However, the target domain may not always have corresponding ready-made domain-specific instruction datasets. Even if they exist, these datasets are often limited in scale.
Several studies investigate domain-specific instruction augmentation, which can be categorized into three aspects: 1) Reusing human-curated public datasets Wang et al. [2023a]; Zhang et al. [2023e]; Jiao et al. [2023]; Lee et al. [2024]; Xia et al. [2024]. For instance, Parrot Jiao et al. [2023] enhances translation capabilities of LLMs in chat by converting bilingual sentences into instruction-following datasets. Furthermore, works like INSTA Lee et al. [2024] and LESS Xia et al. [2024] attempt efficient domain-specific instruction augmentation via dense retrieval. INSTA Lee et al. [2024] uses instructions without responses for effective retrieval. LESS Xia et al. [2024] assumes access to a validation set and uses the warmup LLM’s gradients of the train and validation sets for retrieval. 2) Using seed tasks for self-instruct Luo et al. [2024]; Wan et al. [2023]: Explore-Instruct Wan et al. [2023], for example, employs activate exploration to tree-model domain tasks from both depth and breadth, increasing seed instructions’ coverage of the domain, and subsequently generating broader in-domain instructions through self-instruct. 3) Scaling instructions from the web: Recent works Yue et al. [2024]; Zhou et al. [2024] highlight the immense potential of mining naturally occurring instructions from the internet. Compared to generated data, web-mined instructions exhibit less bias and greater diversity. MAmmoTH2 Yue et al. [2024] firstly retrieves domain-relevant texts and employs a LLM to extract Q-A pairs, further refining them into instruction-response pairs. Jiuzhang3.0 Zhou et al. [2024] distills GPT’s instruction generation capabilities into a smaller model and then uses it to generate instructions from the internet. In conclusion, models fine-tuned with augmented instructions have shown promising domain-specific capabilities.
Appendix C Experiments
C.1 Implementation Details
We consider FedDIT in the cross-device scenario, $N$ = 10 clients, $\mathcal{R}$ = 30 rounds, where we randomly sample 2 clients to be available for each round. Then, each available client performs FedDIT for 10 steps with AdamW optimizer, and the batch size is $B=32$ in a round. The initial learning rate is $55e-$ 5 with a cosine learning rate scheduler. Our experiment utilizes the widely used LLM, Llama3-8B [^2] https://huggingface.co/meta-llama/Meta-Llama-3-8B as the base model with 2048 max sequence length and adopts LoRA tuning method. The rank of LoRA is 16, and the scalar alpha is 16. For k-means Wu [2012], we set cluster num $\xi$ = 10 and for FedDCA we set the similarity threshold $\alpha$ = 0.7. For FedDCA, we set the temperature parameter $\tau$ = 0.5 for contrastive learning Khosla et al. [2020]. We utilize bge-large-en-v1.5333https://huggingface.co/BAAI/bge-large-en-v1.5 as both the client and server’s encoder as default, which outputs embeddings of 1024 dimensions. While we utilize all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 as the client’s small encoder which outputs embeddings of 384 dimensions and bge-large-en-v1.5 as the server’s encoder for FedDCA.
Projector.
@@ C.1 Implementation Details
Prompts Used in the Self-Instruct Data Generation.
@@
Prompt Used for Memory Extraction Attack.
@@
C.2 Baselines
Method Privacy Preserving API Cost Additional Information Performance Degradation Domain Coverage Oriented
FedAvg McMahan et al. [2017]
Random Sampling
LESS Xia et al. [2024]
Self-Instruct Wang et al. [2022]
Direct Retrieval
FedDCA (ours)
Table 6: Key differences between FedDCA and other baselines. FedDCA demonstrates the ability of: 1) privacy-preserving, 2) no API cost, 3) no additional information required, 4) avoiding performance degradation, and 5) aiming at domain coverage optimization. LESS requires additional information, as it needs access to the validation set for gradient-based retrieval.
In Table 6, we highlight five critical aspects of our approach compared with other methods Xia et al. [2024]; Wang et al. [2022]; Zhang et al. [2024b]. Our comparison focuses on five key dimensions:
  • Privacy Preserving: Most methods except Self-Instruct maintain data privacy, we analyze the privacy-preserving ability of FedDCA in Sections 5 and C.3.
  • API Cost: FedDCA requires no api calls. As Self-Instruct utilizes the GPT to generate instructions, it requires additional API calls, which incurs costs.
  • Additional Information: LESS needs validation set access for gradient-based retrieval, whereas FedDCA works with only the basic client’s local training data.
  • Performance Degradation: From Table 2 we see that there occurs performance degradation for FedAvg, Random Sampling, and Self-Instruct, while FedDCA avoids this issue.
  • Domain Coverage: Uniquely, FedDCA is specifically designed to optimize domain coverage for distributed environments, a feature absent in other baseline methods.
Specifically, Self-Instruct Wang et al. [2022] represents the upper limit performance of the generation-based method, which is restricted to the high expense of cost, limited API call frequency, and heavy quality screening process. On the other hand, FewFedPIT Zhang et al. [2024b] attempts to augment local data by leveraging the pre-trained Llama2 model to perform self-instructed generation during training. However, two concerns exist: 1) The pre-trained LLM cannot provide effective, stable, and high-quality instruction generation. 2) Generating instructions locally is very costly regarding both time and computational resources. Overall, generating instructions self-instructively is not a satisfactory method for the FedDIT scenario. LESS Xia et al. [2024] represents the augmentation methods through gradient feature retrieval. However, its performance is underwhelming. Compounding this issue, LESS presupposes access to the validation set, which is not always available, and necessitates an initial warmup on the public dataset before calculating gradients for both the public and validation data. This process becomes computationally burdensome with large public datasets.
In conclusion, these comparisons demonstrate FedDCA’s distinctive approach to balancing privacy preservation, computational overhead and domain coverage optimization, while maintaining minimal additional requirements.
C.3 Domain Inference Attacks
The server may inference the clients’ data domain when the domain data retrieval is performed on the server side. However, the proposed algorithm FedDCA is independent of the presence of a public dataset on the server. Even if the server does not have a public dataset, clients can upload their cluster centers to the server, which selects a set of client centers and sends them back to the clients. Each client can then retrieve data from the website based on the received client center by itself, thereby achieving data augmentation while maximizing the cross-client domain coverage.
In that case, since the server does not know the encoder used by the client, it cannot infer the semantic meaning of the embedding. Thus, for the server, it becomes significantly more challenging to infer the client’s domain, let alone apply any privacy protection techniques to the embeddings.
We provide two examples for illustration. Two different encoders are used as client’s and server’s respectively: BAAI/bge-large-en-v1.5 (denoted as $w_{1}$) and google-bert/bert-large-uncased (denoted as $w_{2}$). Both encoders output 1024-dimensional features.
Example 1: Both $w_{1}$ and $w_{2}$ take "hello world" as input, and the cosine similarity between their embeddings is 0.1829.
Example 2: Three instructions are used:
  • Instruction 1: Create an array of length 5 which contains all even numbers between 1 and 10.
  • Instruction 2: Write a replace method for a string class which replaces the given string with a given set of characters.
  • Instruction 3: What is the sentiment of this news? Please choose an answer from {negative/neutral/positive}. Teollisuuden Voima Oyj, the Finnish utility known as TVO, said it shortlisted Mitsubishi Heavy’s EU-APWR model along with reactors from Areva, Toshiba Corp., GE Hitachi Nuclear Energy, and Korea Hydro & Nuclear Power Co.
For these instructions, Instruction 1 is passed to $w_{1}$ and Instructions 2 and 3 are passed to $w_{2}$, which will result in three embeddings: $e_{1}$, $e_{2}$, and $e_{3}$. The cosine similarity between $e_{1}$ and $e_{2}$ is 0.1464, while the similarity between $e_{1}$ and $e_{3}$ is 0.1879. Instructions 1 and 2 are in the same domain, whereas they have a lower cosine similarity.
In conclusion, as demonstrated above, when the clients do not perform domain-specific instruction retrieval on the server side, the server cannot infer the client’s domain based on the uploaded embeddings.
C.4 How Far can FedDCA Go
Assume an extreme setting, where exists a pre-trained model that has already fine-tuned on the whole in-domain data of the internet or even all open-source public instruction datasets (it is actually unrealistic or temporary, as new data is constantly being generated), is there any room for instruction augmentation or FedDCA? To answer this question, we conduct the following experiments on four settings: A) Fine-tuning on the whole public dataset. B) Following A, perform further federated fine-tuning with only private clients’ local dataset. C) Following A, each client randomly samples 100 public data for fine-tuning. D) Following A, each client samples 100 public data through FedDCA and then performs FedDIT.
Setting H-Eval M-Med FPB FiQA TFNS GSM8K
A 38.41 69.10 75.33 41.09 71.60 53.75
B 43.90 64.30 77.97 66.54 78.26 56.40
C 41.46 66.90 74.50 29.81 70.51 57.01
D 44.12 70.90 81.43 72.00 78.81 58.90
Table 7: Performance (%) of setting A, B, C, and D in each domain.
As shown in Table 7, even with only 100 samples selected via FedDCA, the method not only prevents performance degradation but also helps the model achieve better generalization within the domain. This is attributed to exposure to more diverse in-domain data during training. Meanwhile, it can be observed that random sampling still leads to performance degradation compared to Setting A, and even performs worse than Setting B, where no data augmentation is applied. In the financial domain, performance degradation is also evident.
In conclusion, this experiment further highlights the necessity and effectiveness of FedDCA.
C.5 Augmentation Strategy Visualization
@@
To more intuitively compare the domain coverage of different instruction augmentation methods, we randomly sample 5,000 instructions obtained through these methods and 10,000 in-domain instructions as the background, representing the distribution of specific domains in the public dataset. We then visualized the results using t-SNE van der Maaten and Hinton [2008], as shown in Figure 10. The plot shows that FedDCA encompasses most of the in-domain data, which is consistent with FedDCA’s domain coverage of each domain shown in Table 2. Also, we can observe that the random sampling strategy selects a lot of out-of-domain data while does not have good coverage in specific domains.
(a) Code
(b) Medical
(c) Financial
(d) Mathematical
Figure 10: Visualization of cross-client data distribution in different domains, performing t-SNE dimensionality reduction on retrieved instructions through various augmentation strategies. We randomly sample 10,000 in-domain samples as background while randomly sampling 5,000 samples from the cross-client augmented dataset for different instruction augmentation methods for comparison.